America West Airlines reports the flight time from Los Angeles International Airport to Las Vegas is 1 hour and 5 minutes, or 65 minutes. Suppose the actual flying time is uniformly distributed between 55 and 75 minutes.

What is the probability the flight time is MORE than 68 minutes? Answer to the nearest hundredth (0.01).

My professor uses Microsoft Excel heavily so any commands you may know to help solve the problem would be greatly appreciated!
Answer #1

For this basically we need to know the formula, then we can perform calculation in excel.

It says uniformly distributed between 55 and 75.

And we need to find probability of more than 68 minutes. So, it will take from 68 minutes to 75 minutes (as 75 is max possible value)

\tiny P(x>68)=\frac{75-68}{75-55}=\frac{7}{20}=0.35

So, 0.35 is the probability that it will take more than 68 minutes.

In Excel:

f =(75-68)/(75-55) B1 A 0.35 2 3 4 LO

So, basic need is to understand how to do.

Please comment if any doubt. Thank you.